OpenStudy (anonymous):

use discriminant to determine the number and type of solutions of the equation: 4x^2-12x=-9 choose correct statement: a. the equation has one real number solution b. the equation has two complex number solutions c. the equation has two real number solutions

OpenStudy (anonymous):

first write \[4x^2-12x+9=0\] then compute \[b^2-4ac\] with \[a=4,b=-12,c=9\]you get \[(-12)^2-4\times 4\times 9=0\] which means there is one real solution

OpenStudy (anonymous):

it also means \[4x^2-12x+9\] is a perfect square. it is \[(2x-3)^2\]